# Transfer Learning with Foundational Models for Time Series Forecasting using Low-Rank Adaptations

### v1

@@ Abstract
High computational power and the availability of large datasets have supported the development of Foundational Models. They are a new emerging technique widely used in Generative Artificial Intelligence, characterized by their scalability and their use in Transfer Learning. The enormous and heterogeneous amounts of data used in their initial training phase, known as pre-training, give them a higher generalization capacity than any other specific model, constituting a solid base that can be adapted or adjusted to a wide range of tasks, increasing their applicability. This study proposes LLIAM, the Llama Lora-Integrated Autorregresive Model. Low-Rank Adaptations are used to enhance the knowledge of the model with diverse time series datasets, known as the fine-tuning phase. To illustrate the capabilities of our proposal, two sets of experiments have been carried out that obtained favorable and promising results with lower training times than other Deep Learning approaches. With this work, we also encourage the use of available resources (such as these pre-trained models) to avoid unnecessary and costly training, narrowing the gap between the goals of traditional Artificial Intelligence and those specified by the definition of Green Artificial Intelligence.
@@ 1 Introduction
The advent of Foundation Models (FMs) has led to an explosion in several areas within the Machine Learning (ML) paradigm, and in particular Deep Learning (DL). FMs are a novel type of models that are trained on a large amount of data from different sources, which they use to develop a high generalization capacity. Several fields such as Natural Language Processing (NLP) or Computer Vision (CV) [1] have experienced a great evolution due to the good performance of these models in tasks related to the generation of new and meaningful content from training data, known as Generative AI [2]. In addition, their large generalization capacities allow them to be adapted easily and quickly to tasks from unknown domains. This is known as Transfer Learning (TL) [3].
@@
Due to the great interest in FMs and the opportunities they offer in the field of TSF, we present a method that takes advantage of TL and exploits the capabilities of a LLM to perform this task. This proposal is notable for the fine-tuning of a well-known pretrained FM for language modelling called LLaMA [6] using an efficient technique known as Low-Rank Adaptation (LoRA) [7]. A modification of a prompting technique employed for feeding time-series as textual prompts inside these models is also introduced. The combination of these techniques altogether is called the Llama Lora-Integrated Autorregresive Model (LLIAM).
To demonstrate the capabilities of our proposal respecting different conventional DL methods and its base model (LLaMA) two experiments have been designed and performed, obtaining promising results in both of them. In the first one, we evaluate how LLIAM performs after fine-tuning it over several datasets and compare it to other DL approaches used on TSF. In the second one, we measure whether the previous adjustment improves the results with respect to LLaMA when making predictions on unknown data sets, which is called zero-shot forecasting.
The use of these models, which undergo intensive pre-training, can facilitate the development of new sustainable and environmentally friendly models, aligning with the objectives of Green AI [8]. This approach reduces the need for extensive training to address issues such as hyperparameter optimization or the selection of an ideal architecture. As a result, it allows progress to be made on a variety of topics while maintaining a balance between efficiency and effectiveness.
This paper is structured as follows: Section 2 describes the work related to our proposal, starting with the concept of time series, the main DL methods used for TSF and ending introducing the notion of Foundational Model (FM) and Large Languages Model (LLM). Section 3 presents our proposal, LLIAM, an adaptation of the famous LLaMA efficiently adapted to time series forecasting tasks. Section 4 presents the experimental framework, detailing the materials and methods used for the two experiments performed and how each model is evaluated. The results are also presented in this section. Finally, Section 5 concludes the article with some remarks drawn from the experiments conducted.
@@ 2 Related Work
Throughout the last years, TSF has been of interest to the scientific community. There are a wide number of methods designed to perform this task. In Section 2.1, we discuss the definition of the time series prediction problem, and in Section 2.2 how it has been approached using DL models for sequence modeling. Subsequently, the definition and importance of the Transfer Learning (TL) paradigm applied to Deep Learning (DL) methods is discussed in Section 2.3. Finally, the arrival of FMs and Transformer-based LLMs related to TL is presented in sections 2.4 and 2.5.
2.1 Time-series forecasting
@@
In our day-to-day life, there are numerous phenomena that exhibit different behaviors depending on the time at which they are observed. The measurement of the air temperature at a weather station, the occupancy of any type of service or the stock of specific products are examples whose activity is influenced by the presence of strong or weak time dependencies. Following the example of temperature, it is logical that at noon it is higher than in the morning or at night. Moreover, depending on the season, we can assume its variation.
The main purpose of TSF is to predict the future values of time-series, taking advantage of their temporal dependencies. In essence, this task is described as finding a function $f$ capable of generating $h$ predictions $\hat{Y}=\{\hat{y}_{t+1},\hat{y}_{t+2},...,\hat{y}_{t+h}\}$ given $n$ previous observations $S=\{x_{0},x_{1},x_{2},...,x_{n}\}$, minimizing the error between them and the real $Y=\{y_{t+1},y_{t+2},...,y_{t+h}\}$ ones. In this context, the number of time-steps to predict is called forecast horizon and the historic previous values of a series are called lags.
The pioneers working in this task were primarily statisticians whose algorithms depended on many time-series analysis techniques, expending much of their time designing them to extract and work with different features such as its trend, seasonal patterns or stationality properties.
2.2 Deep Learning and TSF
Machine Learning (ML) [11] is a subset of AI whose objective is bringing the capacity of learning from experience to computers without being explicitly programmed to do so. Inside this, another subset is found, known as DL [12]. Unlike ML methods, DL ones integrate the learning of representation into them, enabling the generation of representations with different levels of abstraction by stacking several processing or hidden layers. Artificial Neural Networks (ANNs) are the fundamental models used in DL techniques, being the artificial neuron the base element of their layers. ANNs are also considered to be universal approximators [13]. However, not all ANNs are considered Deep Neural Networks (DNN). When the number of hidden layers is greater than one, the model is not considered deep. These types of networks are known as Shallow Neural Networks (SNN) [14].
The great performance of DL in various fields [15] resulted in numerous proposals capable of modeling time-series without investing much time performing feature-engineering over them. Sequence-to-sequence (seq-to-seq) architectures are chosen for tasks in charge of generating a sequence through another sequence [16]. Encoder-decoder one is the standard used. It employs an encoder ($E$) for generating a latent representation (or context vector, $z$) of the input sequence, summarizing important information extracted from the series into a set of abstract features. This representation is then used by a decoder ($D$) module for generating the desired output. $E$ and $D$ are usually two DL architectures [17]. Nevertheless, the use of a full encoder-decoder architecture is not mandatory. The following paragraphs introduce the main models proposed for sequential data.
@@ 2.2 Deep Learning and TSF
Recurrent Neural Networks (RNNs) are employed to detect patterns in sequential data such as text, time series or even genomes [18]. RNNs process an input sequence step-by-step, updating a hidden state at each one. This hidden state retains historical information of the sequence up to the current time step, providing an abstract and summarized representation of the entire input sequence once the last step is processed.
Following the seq-to-seq architecture, RNNs can be used as both encoders and decoders [19]. In instances where both components are employed, the encoder RNN processes the input sequence and generates a context vector that is equivalent to its hidden state. Subsequently, the output of the encoder RNN is employed as input for the decoder RNN, which generates the subsequent desired steps in an autoregressive fashion. Another common approach is to replace the decoder RNN with a unique dense layer or a multi-layer perceptron (MLP), which allows the inference of all the required steps in a single forward pass. This strategy is known as Multi-Input Multi-Output (MIMO) and is capable of generating a forecast horizon in a single step [20].
However, it is necessary to address a set of problems concerning these networks when processing long enough series. The loss of long-range dependencies, the vanishing or exploding gradient problem and their sequential nature are the principal ones. Improvements on RNNs have been proposed for mitigating the first two.
Long-Short Term Memory (LSTM) [21] introduce a new state, referred to as long term memory, and three gated mechanisms for updating both states based on the input data and generating the output as a combination of these states. In contrast, Gated Recurrent Units (GRU) [22] do not introduce a second state. Two alternate gates dictates what to forget and introduce into the hidden state, which is used for calculating the output of a cell. Both LSTM and GRU enhance the capacity of RNN for handling long-term dependencies while trying to mitigate the vanishing or exploding gradient problem.
Additional improvements to these networks try to enhance long-term dependencies. Bidirectional recurrent networks apply two LSTMs to the input sequence in both ways, forward and backward [23], and alternative LSTM cells with peephole connections [24] have been proven to enhance LSTM baseline performance when connecting the long-term memory with every gate.
@@
Other proposals also aim to address the issue of computing input series sequentially. Convolutional Neural Networks (CNNs) [25, 26] are characterized by their good performance on image processing, but they have also been used for sequence modeling tasks. The main feature of these networks is the use of convolutions, as their name suggests. A convolution is an operator capable of extracting features sliding a filter, represented as a matrix, over the input data. When the input data are series, it is usual to rely on unidimensional convolutions (1DCNN). Nevertheless, these are not entirely suitable for sequence modeling tasks. These networks may disregard the temporal order of the data, incorporating subsequent observations when processing earlier ones.
Temporal Convolutional Networks (TCNs) [27] combine some of the best practices of modern convolutional architectures for sequence modeling, claiming to outperform RNNs in a broad collection of related tasks. A modification of the baseline convolutional operator originally used for audio generation [28] is employed, known as causal dilated convolution, inside the main component of these networks. This component is called residual block.
The causality of this operator allows the model to respect the temporal order of the data, while the dilatation technique maximizes the coverage of the receptive field for perceiving longer time dependencies without increasing the computational cost or the number of trainable parameters.
@@
Figure 1: Transformer baseline architecture. Adapted from [17].
Transformers [17] arose as state-of-the-art architectures for sequence translation with multiple stacked encoders and decoders, capable of processing the input sequence parallelly and demonstrating superior performance compared to other classic models. As seen in figure 1, encoders and decoders are composed of Multi-Head Attention (MHA) mechanisms and Feed Forward Networks (FFN) with residual connections and layer normalization. It is important to note that while the encoder stack processes all elements of the input series in a single forward pass, the decoder follows an autoregressive fashion when predicting the output sequence.
Three attention procedures can be found in the base Transformer. The one used in the encoder is known as self-attention, and lets the model discover the different relationships between the elements of the input series. In the decoder, masked-attention utilizes a mask to prevent a position from attending future ones. Cross-attention uses the encoders stack output to find the relationships between them and the decoder input. MHA enhances the capacity of the model by simultaneously computing various attentions, referred to as “heads”, which lets it focus on different aspects of the data in each one. Note that the parallel processing of all tokens entails the need to introduce positional information into their representations to ensure that it takes into account the order of the tokens.
While MHA processes relationships between positions, FFNs are implemented at the end of all encoders and decoders to introduce non-linearity in each one separately. Between the MHA and FFN modules, residual connections and a layer normalization are applied to stabilize the learning phase of the model and improve its performance. Residual connections guarantee that the processed representations of the input tokens really represent them, and layer normalization mitigates the vanishing/exploding gradient problem for effective training.
The base architecture of Transformers has been adapted for numerous tasks, including TSF. These modifications are highly relevant today due to their strong performance across various applications. The adaptations range from simple single-component changes to entirely new architectures [29].
@@ 2.3 Transfer Learning
Using DL methods requires abundant amounts of training and test data from a specific domain to resolve tasks. However, in real-world scenarios, obtaining such data can be challenging or extremely expensive. Therefore, it is worth exploring whether knowledge from a known domain can be applied to an unknown one like humans do. This is the main purpose of transfer learning (TL) [30].
@@
TL has been combined with numerous models from different learning paradigms to resolve specif tasks. Some examples follow. Medical images analysis tasks have benefited from this paradigm using pretrained and/or fine-tuned deep classification and segmentation models, coping with the limited amount of labeled data that can be provided in these domains [31]. Reinforcement Learning (RL) techniques have been adapted to integrate knowledge from diverse domains and their tasks, such as game playing, Natural Language Processing (NLP) and training large state-of-the-art models [32].
2.4 Foundation Models
The concept of FM was first introduced by researchers at Stanford University. FMs benefits from TL and scalability, enabled by advances in computing power and the availability of large and diverse datasets. These models serve as a robust base that can be adapted or fine-tuned for a wide range of tasks, enhancing their applicability across various domains [33]. FMs have a remarkable generalization capacity, which allows them to be proficient in several tasks without the need to be shown additional examples (zero-shot learning) or with only a few examples (few-shot learning), because they are able to learn from the context that defines their input (in-context learning).
Moreover, this input can be refined to maximize its performance with prompt-engineering techniques. It is possible to adapt its knowledge to any domain using different Parameter-Efficient Fine-Tuning (PEFT) [34, 35] techniques. This allows an efficient and fast fine-tuning of a model without the need to modify all of its weights. Although these methods arose mainly for the adaptation of LLMs, they are not limited to them or to the Transformer architecture. These can be grouped into four categories [35]: additive, which modify the model architecture to introduce new trainable components; selective, which refine only a subset of the model weights; re-parameterization, which transform the model in its fitting stage, but then integrate these transformations into its original architecture; and hybrid, which combine several methods.
Among these techniques, adapters [36] are a prominent set of methods which introduce small trainable layers between the various modules of the model. Within this category we also find prompt-tuning [37], useful in multitask language models, which seeks to adapt a human input (hard-prompt) to a more affine model (soft-prompt) by operating directly on the continuous vector space given by the model encoder. This overrides the interpretability of the original input. Similar to prompt-tuning are prefix-tuning procedures [38], which only adds a set of trainable vectors for each task faced by the model at the beginning of the input. Other frequently used techniques are Low-Rank Adaptations (LoRA) [7], which fall into the category of re-parameterization. They introduce in parallel in the modules of a model two trainable weight matrices which are combined with each other’s output, and which have a rank much smaller than the dimension of the model to be fitted. These matrices can be integrated into the original model parameters (known as the merge stage) ensuring that their inference time does not vary.
Diverse proposals have exploited the properties of FM for time-series analysis tasks, with Transformer-based architectures beign the most used [39] as they can handle sequential data effectively.
2.5 Large Language Models
LLMs are a specific type of FM trained on large amounts of textual data. Contrary to the pretrained models mentioned in Section 2.3, their superior general-purpose language understanding grants them the ability to resolve various tasks with impressive performance without requiring a fine-tuning phase. Nowadays, most of the actual architectures used by LLMs are based on the Transformer one [17].
The most famous models are currently developed and maintained by large companies. The LLaMA family is a collection of several foundational text models developed in 2013 by Meta, based on a modified Transformer decoder architecture [6]. These models are usually pretrained over a mixture of several publicly-available sources written in different languages, such as Wikipedia, GitHub or ArXiv. Improvements over this family has been made, releasing LLaMA-2 later in the same year [40] and LLaMA-3 in 2024 [41]. The differences between them focus on the increase in the quantity and quality of the data used in their pretraining phase and the use of a larger context window. Another popular family is the Generative Pretrained Transformer (GPT) one, developed and maintained by OpenAI. This was originated as an NLP multitask pretrained decoder-only Transformer model [42] and evolved into a larger and complex multi-modal one in its latest versions, such as GTP-4 [43] and GPT-4o, the last one being capable of processing audio, vision, and text in real time. Not all FMs began as LLMs. The first generation of the Gemini family from Google, for instance, started as a different decoder-only Transformer multimodal architecture. This design is suitable for various applications, from complex reasoning tasks to those requiring on-device memory constraints [44].
@@ 2.5 Large Language Models
Time-series tasks have increasingly benefited from TL. Pretrained models, as discussed in Section 2.1, have been employed for the classification, forecasting, clustering, anomaly detection, and imputation of time-series data [45]. Transformer-based architectures are particularly significant, serving as the foundation for many models, including the well-known LLMs. These models can abstract various language domains due to the vast amounts of diverse data they are trained on. This generalization ability suggests potential applications beyond language tasks, which will be applied next, such as zero-shot TS predictors [5].
Figure 2: LLIAM arquitecture.
3 LLIAM: Llama Lora-Integrated Autorregresive Model

### v3

@@ Abstract
Foundational Models are an emerging widely used technique of Generative Artificial Intelligence. These models are distinguished by their scalability and the ease with which they can be adapted through the exploitation of Transfer Learning. The availability of high computational power and large datasets have supported their development, achieving a high generalization capacity due to the enormous and heterogeneous amounts of data used in their initial training. These characteristics contribute to a solid base that can be adapted or adjusted to a wide range of tasks, increasing their applicability. This study proposes the methodology LLIAM, a straightforward adaptation of a kind of Foundational Models, Large Language Models, for the Time Series Forecasting task. An adequate time-series prompting schema and Low-Rank Adaptations are used to enhance the knowledge of the model with diverse time series datasets, known as the fine-tuning phase. A study divided in two stages has been performed for evaluating the effectiveness of the proposed methodology. Initially, a comparison was made between the performance of LLIAM and different state-of-the-art Deep Learning algorithms, including Recurrent Neural Networks and Temporal Convolutional Networks, as well as a LLM-based method, TimeLLM. Following this, a zero-shot study is presented in order to evaluate the generalization capacity of the proposed methodology with time series datasets from unknown domains not considered in the model training. The outcomes of this investigation demonstrate the efficacy of LLIAM, highlighting that this straightforward and general approach can attain competent results without the necessity for applying complex modifications. This work also encourages the use of available resources (such as these pre-trained models) and efficient fine-tuning techniques to avoid unnecessary and costly training, narrowing the gap between the goals of traditional Artificial Intelligence and Green Artificial Intelligence.
@@ 1 Introduction
The advent of Foundation Models (FMs) has set a milestone in several areas within the Machine Learning (ML) paradigm, and in particular Deep Learning (DL). FMs are a type of models that are trained on a large amount of data from different sources, which they use to reach a high generalization capacity integrating all the knowledge extracted form these sources into them. Several fields such as Natural Language Processing (NLP) or Computer Vision (CV) [1] have experienced a great evolution due to the good performance of these models in tasks related to the generation of new and meaningful content from training data, known as Generative AI [2]. In addition, their large generalization capacities allow them to be adapted easily and quickly to tasks from unknown domains through Transfer Learning [3].
@@
Due to the great interest in FMs, especially LLMs, and the opportunities they offer in the field of TSF, we were interested in whether: i) can we take advantage of the similarities between text and time-series and the high abstraction capacity of LLMs for the time series task? ii) is a large adaptation of these models necessary to achieve good results? It is for these reasons that we present a methodology that takes advantage of Transfer Learning and exploits the capabilities of a LLM to perform this task. This proposal is notable for the fine-tuning of a well-known pre-trained FM for language modeling called LLaMA [6] using an efficient technique known as Low-Rank Adaptation (LoRA) [7]. A modification of a prompting technique employed for feeding time-series as textual prompts inside these models is also introduced. The proposal, based on the combination of these techniques altogether is called the Llama Lora-Integrated Autoregressive Model (LLIAM).
An experimental study with two groups of experiments has been designed to demonstrate the capabilities of our proposal respecting different conventional DL methods, and LLM-base approaches: its base model (LLaMA) and a state-of-the-art model, TimeLLM. In the first one, we evaluate how LLIAM performs after fine-tuning it over several datasets and compare it to other approaches used on TSF. In the second one, we measure the behaviour of LLIAM when making predictions on data sets from unknown domains, which is called zero-shot forecasting.
The use of approaches as LLIAM, which undergo intensive pre-training, can facilitate the development of new sustainable and environmentally friendly models, aligning with the objectives of Green AI [8]. LLIAM reduces the need for extensive training to address issues such as hyperparameter optimization or the selection of an ideal architecture.
This paper is structured as follows: Section 2 describes the work related to our proposal, starting with an introduction to the TSF task and the main DL methods that have been used along time. Next, a description of Transfer Learning, Pre-trained Model, FM and LLM with proposals that have benefited from their properties applied to TSF is presented. Section 3 presents our proposal, LLIAM, an adaptation of the famous LLaMA efficiently adapted to the TSF task. Section 4 presents the experimental framework, detailing the materials and methods used for the two experiments performed and how each model is evaluated. The results are also presented in this section. Finally, Section 5 concludes the article with some remarks drawn from the experiments conducted.
@@ 2 Related Work
Throughout the last years, a wide number of methods have been designed to perform time-series analysis tasks, including TSF. In Section 2.1, the fundamental concepts for understanding the TSF problem and a brief explanation of the main DL methods that have been used are introduced. After that, the definition of Transfer Learning is discussed in Section 2.2 with special emphasis on the definitions of Pretrained Models, FM and LLM and how their propoerties have caused a paradigm-shift in the resolution of the TSF task. Finally, in Section 2.3 is presented some of the most relevant tuning techniques that allow these models to be adapted to a wide range of tasks.
2.1 Temporal prediction
In our day-to-day life, there are numerous phenomena that exhibit different behaviors depending on the time at which they are observed. The measurement of the air temperature at a weather station, the occupancy of any type of service or the stock of specific products are examples whose activity is influenced by the presence of strong or weak time dependencies. Following the example of temperature, it is logical that at noon it is higher than in the morning or at night. Moreover, depending on the season, its variation may be assumed. This type of data which is dependent on time is called time-series.
@@ 2.1 Temporal prediction
The resolution of the TSF task has always been an interesting and challenging one due to the wide range of domains that present time-dependent information. It is not easy to model the temporal dependencies contained in time series data, as how these will develop. In essence, this task is described as finding a function $f$ capable of generating $h$ predictions $\hat{Y}=\{\hat{y}_{t+1},\hat{y}_{t+2},...,\hat{y}_{t+h}\}$ given $n$ previous observations $S=\{x_{0},x_{1},x_{2},...,x_{n}\}$, minimizing the error between them and the real $Y=\{y_{t+1},y_{t+2},...,y_{t+h}\}$ ones. In this context, the number of time-steps to predict is called forecast horizon and the historic previous values of a series are called lags.
Machine Learning (ML) techniques show considerable potential for solving this task. ML is a subset of Artificial Intelligence that facilitates the capacity for computers to learn from experience and automatically extract patterns from data. However, it should be noted that the implementation of these techniques necessitates a preliminary investment in the curation and enhancement of the quality of the features presented in the data, including the creation of new ones. A subset of ML, called Deep Learning (DL), aims to minimize this effort by integrating the learning of representation into them, stacking several processing or hidden layers responsible for generating representations of the data with different abstraction levels.
Artificial Neural Networks (ANN) are the base models employed in DL, with the artificial neuron constituting the base element of their layers. ANNs are considered universal approximators [11]. However, it should be noted that not all ANNs are considered to be deep (Deep Neural Networks, DNNs). Simple ANNs with one hidden layer are usually considered Shallow Neural Networks (SNNs) [12].
The great performance of DL in various fields [13] resulted in numerous proposals capable of modeling time-series without investing much time performing feature-engineering over them. Sequence-to-sequence (seq-to-seq) architectures are usually chosen for tasks in charge of generating a sequence through another sequence [14]. Encoder-decoder one is the standard used. It employs an encoder ($E$) for generating a latent representation (or context vector, $z$) of the input sequence, summarizing important information extracted from the series into a set of abstract features. This representation is then used by a decoder ($D$) module for generating the desired output. $E$ and $D$ are usually two DL architectures [15]. Nevertheless, the use of a full encoder-decoder architecture is not mandatory. The following paragraphs introduce the main models proposed for sequential data.
@@
Recurrent Neural Networks (RNNs) are employed to detect patterns in sequential data such as text, time series or even genomes [16]. RNNs process an input sequence step-by-step, updating a hidden state at each one. This hidden state retains historical information of the sequence up to the current time step, providing an abstract and summarized representation of the entire input sequence once the last step is processed. It is necessary to address a set of problems concerning these networks when processing long enough series. The loss of long-range dependencies, the vanishing or exploding gradient problem and their sequential nature are the principal ones. Improvements on RNNs have been proposed for mitigating the first two, such as Long Short Term Memory (LSTM) [17] cells or Gated Recurrent Units (GRU) [18] .
@@
Other proposals also aim to address the issue of computing input series sequentially. Convolutional Neural Networks (CNNs) [19, 20] are characterized by their good performance on image processing, but they have also been used for sequence modeling tasks. The main feature of these networks is the use of convolutions, as their name suggests. Nevertheless, these are not entirely suitable for sequence modeling tasks. These networks may disregard the temporal order of the data, incorporating subsequent observations when processing earlier ones. Temporal Convolutional Networks (TCNs) [21] combine some of the best practices of modern convolutional architectures for sequence modeling, claiming to outperform RNNs in a broad collection of related tasks. A modification of the baseline convolutional operator originally used for audio generation [22] is employed, known as causal dilated convolution, inside the main component of these networks. This component is called residual block. The causality of this operator allows the model to respect the temporal order of the data, while the dilatation technique maximizes the coverage of the receptive field for perceiving longer time dependencies without increasing the computational cost or the number of trainable parameters.
Figure 1: Detailed baseline Transformer architecture. Encoders capture the relationships between input tokens using multi-head attention and enhance their representations with feed-forward layers. Decoders relate the output of the encoder stack to the predicted tokens using masked and cross-attention mechanisms. Both use positional encoding. Post-processing phase (red block) is usually used to predict the next token according to the task using the output of the decoder stack. Figure adapted from [15]
@@
Transformers [15] arose as state-of-the-art architectures for sequence translation with multiple stacked encoders and decoders, as illustrated in Figure 1, capable of processing the input sequence parallelly and demonstrating superior performance compared to other classic models. Encoders and decoders are composed of Multi-Head Attention (MHA) mechanisms and Feed Forward Networks (FFN) with residual connections and layer normalization. It is important to note that while the encoder stack processes all elements of the input series in a single forward pass, the decoder follows an autoregressive fashion when predicting the output sequence. MHA enhances the capacity of the model by simultaneously computing various attentions, referred to as “heads”, which lets it focus on different aspects of the data in each one. Note that the parallel processing of all tokens entails the need to introduce positional information into their representations to ensure that it takes into account the order of the tokens. The base architecture of Transformers has been adapted for numerous tasks, including TSF. These modifications are highly relevant today due to their strong performance across various applications. The adaptations range from simple single-component changes to entirely new architectures [23]. Diverse proposals have leveraged Transformer architectures as the primary framework for TSF tasks. Informer [24] was one of the first adaptations of this architecture, addressing key issues found in the standard model by introducing a novel attention mechanism that reduces time and memory complexities and using a distillation operation to remove redundant information. DSFormer [25] employs attention mechanisms to fuse information from different levels, effectively capturing significant patterns in both the temporal and feature dimensions of the data, alongside specialized sampling methods. PatchTST [26] handles multivariate time-series by isolating features into separate channels that are divided into patches for either supervised or self-supervised training. These patches are then processed by a Transformer backbone to generate predictions.
@@ 2.2 Transfer Learning
Using DL methods requires abundant amounts of training and test data from a specific domain to resolve tasks. However, in real-world scenarios, obtaining such data can be challenging or extremely expensive. Therefore, it is worth exploring whether knowledge from a known domain can be applied to an unknown one like humans do. This is the main purpose of Transfer Learning [27].
@@
Originally, models were pre-trained on large datasets for specific time series tasks like classification, forecasting, clustering, anomaly detection, and imputation [28]. This setup enabled a relatively easy adaptation of models to new domains within the same task category. Such models are known as Pre-trained Models.
@@
FM are an emergent set of models which leverages the capabilities of traditional Pretrained Models, being Transfer Learning one of their fundamental characteristic. Nowadays, a type of FMs called LLMs are widely recognized. These models can abstract various language domains due to the vast amounts of diverse data they are trained on. This generalization ability suggests potential applications beyond language tasks, which will be applied next, such as zero-shot time-series predictors [5]. The next paragraphs describe the concept of FM and LLM.
Foundation Model
The idea of FMs originated with researchers at Stanford University. FMs leverage Transfer Learning and scalability, thanks to advancements in computational power and the wealth of large, diverse datasets. These models act as a robust foundation, adaptable and fine-tunable for an array of tasks, thereby increasing their utility across various fields [29]. A notable characteristic is their exceptional generalization ability, which enables them to perform multiple tasks without additional examples (zero-shot learning) or with minimal examples (few-shot learning), as they can infer from the input context (in-context learning). In essence, FMs signify a paradigm shift in the domain, promoting uniformity with a single model’s application to multiple tasks and fostering new functionalities by leveraging their extensive knowledge base.
Large Language Models
LLMs represent a specialised category of FMs trained on extensive text data. Their exceptional general-purpose language comprehension equips them to tackle a range of tasks effectively without the need for further fine-tuning. Currently, most LLM architectures are built on the Transformer model [15]. The most prominent models in this domain are developed and supported by large corporations. The LLaMA series, introduced by Meta in 2013, includes several foundational language models that were crafted using an adapted Transformer decoder architecture [6]. These models typically undergo pre-training on a mix of publicly accessible sources in different languages, including Wikipedia, GitHub, and ArXiv. Upgrades to this series have led to the launch of LLaMA-2 later in 2013 [30] and LLaMA-3 in 2024 [31], with distinctions mainly in the scale and quality of pretraining data and an expanded context window. Another notable series is the Generative Pre-trained Transformer (GPT) models, created and maintained by OpenAI. Initially formulated as an NLP multitask pre-trained decoder-only transformer [32], these models have evolved into more extensive and sophisticated multimodal systems in their latest iterations, such as GPT-4 [33] and GPT-4o, the latter being capable of handling audio, visual, and textual data in real time. Not all FMs originated as LLMs. For example, Google’s Gemini family commenced with a unique decoder-only Transformer multimodal architecture. This structure is adaptable for various applications, from complex reasoning tasks to scenarios with limitations in memory on the device [34].
Use of FM for TSF
Due to the similarities between textual and sequential data, various new proposals have arisen to exploit the properties of existing pre-trained and FMs for their adaptation to time-series analysis tasks. Other works opted for the creation of new models using recognized architectures proposing methods to align the representation of time-series with the model original input data. Transformer-based architectures are heavily utilized as the main backbone of these methods [35] as they can handle sequential data effectively, but is not limited to only these. STEP [36] leverages the possibilites of Spatio-Temporal Graph Neural Networks (STGNN) creating a pre-trained time-series model built in a two-stage process. First, a transformer model is pre-trained to generate patches (or segment-level representations) containing relevant contextual information from the input time series using the masked autoenconding strategy. Then, a STGNN leverages these representations to make long-term forecasts of the input series. Other proposals may use LLMs as their base considering the similarities between temporal and textual data. TimeLLM [37] is an LLM reprogramming framework that facilitates the adaptation of these models to the TSF, capable of combining numerical and domain-specific information without requiring their fine-tuning. It introduces two complementary techniques to improve LLMs reasoning capacities related to time-series, Prompt-as-Prefix and Patch Reprogramming. Prompt-as-Prefix enhances the adaptability and guidance of LLMs using a textual description of the dataset domain, statistics and the instruction of the task to be resolved. Patch Reprogramming projects the input window into the LLM data representation space, i.e, match time-series to a combination of several sets of textual tokens presented in the model vocabulary (prototypes) such as ”steady down” or ”short up”. According to the authors, this activates the LLM ability to understand and reason with time series data. Chronos [38] is another framework that employs LLM as its backbone using scaling and quantization techniques to tokenize the time-series, being able to adapt these models without changing anything from their design. This straightforward process achieved decent results, although a full training of the model with large amounts of datasets is needed to achieve them. The Frozen Pre-trained Transformer (FPT) architecture [39] originated from a unified framework capable of adapting LLMs to diverse Time-Series Analysis tasks. It freezes the MHA and feed-forward layers of the pre-trained model and fine-tune the rest, reducing the cost of training a complete model. However, the output layer needs to be adapted according to the desired task. Lastly, other proposals focus on adapting the model by attempting to guide the LLM correctly on the basis of its input prompt alone, such as Promptcast [40]. This prompting schema conditions the model to produce a consistent output for the TSF task by introducing information about the data domain, the list of input lags, and a question indicating the number of instants to predict.
2.3 Fine-tuning of LLM
Fine-tuning helps FM achieve optimal performance for specific tasks using small and task-specific data sets. The large number of parameters in these models makes it impractical to modify all of them. Parameter Efficient Fine-Tuning (PEFT) techniques [41, 42] allow for efficient and fast fine-tuning of a model by changing only a small and selected number of weights while leaving the rest unchanged. Although these methods arose mainly for the adaptation of LLMs, they are not limited to them or to the Transformer architecture. These can be grouped into four categories [42]: additive, which modify the model architecture to introduce new trainable components; selective, which refine only a subset of the model weights; re-parameterization, which transform the model in its fitting stage, but then integrate these transformations into its original architecture; and hybrid, which combine several methods.
Adapters
Adapters [43] are a prominent set of methods which introduce small trainable layers between the various modules of the model. Within this category we also find prompt-tuning [44], useful in multitask language models, which seeks to adapt a human input (hard-prompt) to a more affine model (soft-prompt) by operating directly on the continuous vector space given by the model encoder. This overrides the interpretability of the original input. Similar to prompt-tuning are prefix-tuning procedures [45], which only adds a set of trainable vectors for each task faced by the model at the beginning of the input.
Low-Rank Adaptations
Low-Rank Adaptations (LoRA) [7], which fall into the category of re-parameterization. They introduce in parallel in the modules of a model two trainable weight matrices which are combined with each other’s output, and which have a rank much smaller than the dimension of the model to be fitted. These matrices can be integrated into the original model parameters (known as the merge stage) ensuring that their inference time does not vary.
3 LLIAM: Llama Lora-Integrated Autoregressive Model